# Token Communications: A Unified Framework for Cross-modal Context-aware Semantic Communications

### v1

@@ Abstract
In this paper, we introduce token communications (TokCom), a unified framework to leverage cross-modal context information in generative semantic communications (GenSC). TokCom is a new paradigm, motivated by the recent success of generative foundation models and multimodal large language models (GFM/MLLMs), where the communication units are tokens, enabling efficient transformer-based token processing at the transmitter and receiver. In this paper, we introduce the potential opportunities and challenges of leveraging context in GenSC, explore how to integrate GFM/MLLMs-based token processing into semantic communication systems to leverage cross-modal context effectively, present the key principles for efficient TokCom at various layers in future wireless networks. We demonstrate the corresponding TokCom benefits in a GenSC setup for image, leveraging cross-modal context information, which increases the bandwidth efficiency by 70.8% with negligible loss of semantic/perceptual quality. Finally, the potential research directions are identified to facilitate adoption of TokCom in future wireless networks.
@@ I Introduction
Existing wireless networks are mostly designed at the technical communication level, focusing on the reliable transmission of data over noisy channels at high throughput and low latency. This design often disregards the semantic meaning, goal, and context of the transmitted data, leading to communication inefficiencies, especially in unfavorable channel/network conditions and bandwidth-constrained mediums. However, in many emerging applications such as the wireless metaverse and immersive communications, e.g., extended/mixed reality (XR/MR), the internet of senses, and holographic teleportation, effective transmission of the most meaningful and relevant information to the communication goal/intent, is more crucial than character/pixel/sample-exact reconstruction of the source signals. This has led to extensive research on development of efficient semantic communication (SemCom) systems in recent years [1, 2], which are mostly empowered by artificial intelligence and machine learning (AI/ML)-assisted signal processing. More recently, generative AI (GenAI) models, e.g., diffusion models, GANs, and VAEs, have proven to significantly enhance communication at the semantic-level [3, 4, 5]. In generative SemCom (GenSC), the semantics of interest are extracted at the transmitter, communicated over the channel, and then used at the receiver to guide a generative model to locally synthesize a semantically consistent signal with high fidelity. The recent success of powerful generative foundation models (GFMs) and multimodal large language models (MLLMs), e.g., PaLM-E, Sora, and GPT-4o, provides ample opportunities to develop ultra-low rate semantic communication systems. The pre-trained nature of such models and their applicability to a vast range of synthesis tasks, can revolutionize GenSC enabling intent/task-adaptive SemCom systems empowered by pre-trained GFM/MLLMs.
Despite the above advancements, integrating cross-modal context information into semantic communication systems remains less studied. In this paper, we introduce token communications (TokCom), a new GenSC framework that leverages tokens as units of semantic content transmitted within the future wireless networks. Tokens are compressed representations that capture meaning through features of information-rich multimodal data, enabling efficient semantic communications. TokCom is motivated by the recent advent of powerful GFM/MLLMs based on transformer neural networks (NNs), where tokens are the basic processing units of text, images, audio/video signals that may represent words, image patches, temporal audio slices, or video subframes. Tokenization allows complex multimodal data to be broken into manageable, interpretable segments to be processed with state-of-the-art GFM/MLLMs at the TokCom transmitter/receiver, as depicted in Fig. 1. GFMs/MLLMs-based token processing enables to encode the semantic content of multimodal signals at the transmitter and recover the corresponding semantics at the receiver leveraging the cross-modal context information. As an example, a corrupted TokCom packet may lead to an incomplete message like “A beach with palm [MASK] and clear blue water” at the receiver. Here, the missing word is replaced by a [MASK] token. The semantic error correction mechanism in TokCom leverages a pre-trained LLM to predict the masked word based on the surrounding context, outputting, for example, “trees” to complete the sentence as, “A beach with palm trees and clear blue water” thereby avoiding packet re-transmissions that would be required without leveraging context information via the LLM.
@@
  • We introduce a novel TokCom framework along with its basic designs to leverage cross-modal context information in the semantic source compression, semantic channel coding, semantic multiple access, and semantic networking setups. TokCom integrates transformer-based next/masked token prediction via GFM/MLLMs into the transmitter/receiver pipeline to achieve cross-modal context-awareness.
  • We introduce and demonstrate a token-level loss/error mitigation scheme that leverages token likelihood estimation based on cross-modal context to predict and mitigate the tokens corrupted or lost in communication. For a generative image SemCom task, this method improves the bandwidth efficiency by 70.8% at a negligible loss of visual quality by avoiding token retransmissions.
@@ II Tokenization, and Embedding of Various Data Modalities
To process large multimodal data, it is first segmented into various chunks, each consisting of several tokens. Tokens are then assigned unique IDs, and the set of all possible tokens makes a vocabulary. Each token ID is then mapped to a vector that represents the semantic meaning of the token in a dense, fixed-dimensional space, i.e., the embedding space. The embedding vectors are learned via pre-training to capture both syntactic and semantic relationships between tokens. These embeddings allow multimodal language models process inputs more effectively by capturing token semantics as well as the relationships between different tokens, i.e., context. A Tokenizer generates a large codebook of token embeddings that represent the whole corpus of data, e.g., colossal clean crawled corpus (C4) for text, LAION-5B for image-text pairs, and then the (pre-)trained MLLM essentially learns relations between tokens capturing the semantics and context. GFM/MLLMs typically learn a unified token vocabulary valid for all modalities where textual or visual tokens representing similar semantics are mapped to highly correlated vector embeddings. As an example, a vision language model, e.g., the popular contrastive language-image pretraining (CLIP) model [6], learns through pre-training on large corpus of image-text pairs to relate the tokens of a “book” image with its corresponding text token. In the following, we illustrate these concepts for various data modalities.
@@
Text Tokenization/Embedding: In text, a token can represent a word, part of a word, or even a single character, depending on the tokenization model. Commonly, text is split into subwords to handle out-of-vocabulary words. For example, the sentence “The supermarket is hosting a sale” would be tokenized as “-The -super market -is -host ing -a -sale” using WordPiece. Tokenization algorithms like WordPiece, byte-pair encoding (BPE), and unigram language modeling split words into subwords, with each subword mapped to a unique token in the vocabulary. Tokens are then projected into a high-dimensional feature space using word embeddings. Early models like Word2Vec and GloVe used fixed embeddings, while modern transformer models such as bidirectional encoder representations from transformers (BERT) and GPT learn contextual embeddings, meaning the representation of a word like “bank” varies depending on its context. For example, the word “bank” will have different embeddings when used in the context of “a financial institution” versus “the edge of a river”.
@@
Image/Video Tokenization/Embedding: Image tokens are created by dividing the image into fixed-size patches, which are then flattened and embedded. The challenge lies in balancing image fidelity and patch size; smaller patches can increase computational load, while larger ones may lose detail. For instance, vision transformer (ViT) uses $16\times 16$ pixel patches. In addition to patch-based tokenization, vector quantized-variational autoencoder (VQ-VAE) tokenizes images or videos by encoding them into a discrete latent space, where each image is represented by a sequence of tokens from a codebook [7]. This enables more efficient representation and processing by transformer models.
@@
Audio Tokenization/Embedding: Audio signals are typically tokenized by transforming the waveform into spectrograms, typically using techniques such as log Mel filterbanks, which capture both the temporal and frequency-based features of the audio signal. The spectrograms are then divided into overlapping patches, analogous to how images are patchified. Each patch is then flattened and projected into the embedding space to be processed by attention-based transformer models.
@@ III Tokens, Attention, and Transformers
Tokens are commonly used in various vision and natural language processing (NLP) models to represent data in a structured, numerical format that models can process. While transformer models have popularized token-based processing due to their architecture, tokens have long been used in other types of models as well. Transformer-based models, e.g., ViT, BERT, and GPTs, revolutionized token processing by introducing a fundamentally different approach to handling token sequences based on the attention mechanism. Instead of working token by token, the transformer processes the entire sequence of tokens in parallel. This parallelism allows for much faster training and inference, especially on modern hardware like GPUs. Transformers use self-attention mechanisms, where each token directly attends to all other tokens in the sequence, making it much easier for the model to capture long-range dependencies without the constraints of sequential steps. The model dynamically computes attention weights to decide which tokens are important to focus on. This gives transformers greater flexibility in handling long sequences and maintaining more contextual information. Compared to other neural language models, Transformers scale much more effectively [8], making it possible to train very deep models, such as the GPTs, LLaMA, or PaLM with billions of parameters. Transformers are also much easier to train, thanks to their parallelism and efficient gradient flow. They achieve faster convergence and can be trained on much larger datasets. Transformer-based models may learn and predict contextual information either in a unidirectional/auto-regressive or bidirectional fashion.
Unidirectional context learning/prediction refers to models that process input data in a single direction, e.g., left-to-right for text or row-by-row for images, where each token is predicted based only on the preceding tokens. Auto-regressive models like GPTs use this approach and are suited for tasks such as text completion or machine translation, where tokens are generated sequentially without access to future tokens. Fig. 2(a) shows image generation with unidirectional context prediction, where tokens are generated sequentially from a codebook using a unidirectional transformer and then de-tokenized into an image.
Bidirectional context learning/prediction leverages the concept of masking, where random tokens in the input text or image are hidden, and the model learns to predict them by attending to both preceding and succeeding tokens. This bidirectional attention helps capture richer context. Models like BERT use masked language modeling for text, while in visual synthesis, models like MaskGIT [7] predict masked image tokens by considering both nearby and distant tokens, enabling the capture of global spatial dependencies. Fig. 2(b) depicts the typical architecture for bidirectional context learning/prediction. Masked modeling is widely applied to generation tasks, such as image generation [7], where the model iteratively generates image tokens, revealing a portion of tokens at each step and progressively uncovering more with each iteration.
@@
Computational Complexity Typically uses smaller task-specific NN models. Typically based on complex but task-adaptive GFMs/MLLMs, GPTs, etc.
IV Token Communications: Opportunities and Challenges
TokCom is a new GenSC scheme that applies GFMs/MLLMs-based token processing to encode the semantic content of multimodal signals at the transmitter and recover the corresponding semantics at the receiver leveraging cross-modal context information. TokCom provides the following main opportunities/advantages:
  • Tokens unify modalities and TokCom can leverages cross-modal relations to achieve ultra-low-rate SemCom. Furthermore, the in-context learning capabilities [9] of MLLMs enables developing efficient TokCom frameworks that can adapt to various tasks, such as reconstruction, generation, and segmentation, among different modalities.
Table I provides a comparison between TokCom and the conventional SemCom schemes. Despite the above benefits, the large computational complexity of many existing GFMs/MLLMs is a challenge to be addressed in design of efficient TokCom systems. An effective solution to this would be collaborative cloud-edge-device TokCom based on task offloading, which we further investigate in the next sections. It should be noted that the continuous development of smaller models, e.g., small language models (SLMs), and the emergence of more efficient processors will also help overcome this challenge.
@@ V-A TokCom and Semantic Source Compression
TokCom systems can leverage GFMs/MLLMs-based context processing to achieve ultra-low-rate semantic source compression in future wireless networks. Language modeling is inherently a form of compression because it involves accurate prediction the next/masked tokens in a sequence, thereby reducing uncertainty about the data. As the source coding theorem states, the expected length of a source compressed losslessly, is tied to its negative log-likelihood under its probability model. Conventional lossless source coding techniques like arithmetic coding leverage predicted probabilities to encode sequences into a nearly optimal binary form, where more accurate predictions yield shorter codes. State-of-the-art GFM/MLLMs predict the conditional probabilities of tokens given context. The predictability-compression link is general across various data modalities, and by minimizing the log-likelihood loss during training, language models implicitly optimize for compression. Thereby, compression is achieved when models effectively capture patterns and dependencies in the input multimodal data, making GFM/MLLMs inherently capable of functioning as efficient semantic compressors. As an example, the authors in [10] demonstrated that Chinchilla 70B, i.e., a language model primarily trained on text, can efficiently compress ImageNet patches to 43.4% of their original size and LibriSpeech audio samples to 16.4%, outperforming modality-specific compressors like PNG (58.5%) and FLAC (30.3%). Additionally, since tokenizers transform multimodal data into discrete tokens, tokenization inherently acts as a form of lossy compression. Its performance limits are dictated by the rate-distortion-perception theory [11], which defines the trade-off between rate, distortion, and the perceptual quality of the reconstructed signal. Moreover, MLLMs can selectively discard irrelevant tokens based on downstream tasks, enabling efficient lossy token compression. We demonstrate TokCom for semantic source compression in Fig. 3(a).
Figure 3: Basic TokCom setups: (a) Semantic source compression: Left - lossless compression via arithmetic coding and MLLMs for probability prediction; right - lossy compression dropping unimportant tokens based on downstream tasks. (b) Semantic channel coding: Map the embedding space to the modulation space; closer constellations for embedding vectors with high cosine similarity. (c) Semantic multiple access: Tokens from two users overlap in non-orthogonal multiple access, MLLMs separate sources using context and semantic orthogonality. (d) Semantic network protocols: Transmitter packetizes tokens masking some of them for bitrate control; receiver predicts masked/lost packets using MLLMs.
@@ V-B TokCom with Semantic Channel Coding
Another possibility is to leverage the semantic similarity of tokens to optimize a mapping from the token codebook to the channel symbols. For example, the two tokens “King” and “Queen” are semantically very similar. In fact, if the gender information is lost or mistaken due to channel errors, it can be simply inferred or corrected from numerous contextual clues at the receiver leveraging GFM/MLLMs. In other words, semantically similar tokens should be mapped onto channel symbols that are more probable to be confused during transmission. State-of-the-art multimodal embedding techniques enable the semantic similarity of tokens to be measured by their distance in the embedding space [6]. We demonstrate this concept in Fig. 3(b).
@@ VI-A TokCom Setup
Each image is first tokenized into a sequence of $N=256$ discrete tokens. To facilitate efficient transmission, these tokens are grouped into packets, with each packet containing 16 tokens. To mitigate the impact of burst errors—where consecutive erroneous tokens can significantly degrade reconstruction quality due to the loss of contextual information—the token positions included within each packet are randomized. This ensures that adjacent tokens in the original sequence are less likely to be lost simultaneously. The permutation pattern of token indices is pre-determined and remains consistent between the transmitter and receiver. Each token is drawn from a token codebook of size $Q=1024$, meaning it is represented using a 10-bit binary sequence.
For reliable transmission, each packet undergoes rate-$\frac{1}{2}$ convolutional encoding with cyclic redundancy check to enhance error detection and correction. The encoded bits are then modulated using 16-QAM before being transmitted over the channel. On the receiver side, demodulation and decoding are performed using a Soft Viterbi decoder. Based on the packet length, coding, and modulation scheme, the packet error rate (PER) as a function of the signal-to-noise ratio (SNR) under flat Rayleigh fading channel is derived in [13]. Then, the average number of packet retransmissions required for successful data reception, denoted as $T$, is calculated as $\frac{1}{1-\text{PER}}$. Finally, we introduce a key performance metric token communication bandwidth efficiency (TCE), defined as $\frac{h\times w}{T\times N\times\log_{2}({Q})}$.

### v3

# Token Communications: A Large Model-Driven Framework for Cross-modal Context-aware Semantic Communications

@@ Abstract
In this paper, we introduce token communications (TokCom), a large model-driven framework to leverage cross-modal context information in generative semantic communications (GenSC). TokCom is a new paradigm, motivated by the recent success of generative foundation models and multimodal large language models (GFM/MLLMs), where the communication units are tokens, enabling efficient transformer-based token processing at the transmitter and receiver. In this paper, we introduce the potential opportunities and challenges of leveraging context in GenSC, explore how to integrate GFM/MLLMs-based token processing into semantic communication systems to leverage cross-modal context effectively at affordable complexity, present the key principles for efficient TokCom at various layers in future wireless networks. In a typical image semantic communication setup, we demonstrate a significant improvement of the bandwidth efficiency, achieved by TokCom by leveraging the context information among tokens. Finally, the potential research directions are identified to facilitate adoption of TokCom in future wireless networks.
@@ I Introduction
Motivated by emerging applications, recent research has focused on development of efficient semantic communication (SemCom) systems [1, 2], which are mostly empowered by artificial intelligence and machine learning (AI/ML)-assisted signal processing. More recently, generative AI (GenAI) models, e.g., diffusion models, GANs, and VAEs, have proven to significantly enhance communication at the semantic-level through generative SemCom (GenSC) [3, 4]. The recent success of powerful generative foundation models (GFMs) and multimodal large language models (MLLMs), e.g., PaLM-E, Sora, and GPT-4o, provides ample opportunities to develop ultra-low bitrate semantic communication systems. The pre-trained nature of such models and their applicability to a vast range of synthesis tasks, can revolutionize GenSC enabling intent/task-adaptive SemCom systems empowered by pre-trained GFM/MLLMs.
Despite the above advancements, integrating cross-modal context information into semantic communication systems remains less studied. In this paper, we introduce token communications (TokCom), a new GenSC framework that leverages tokens as units of semantic content transmitted within the future wireless networks. Tokens are compressed representations that capture meaning through features of information-rich multimodal data, enabling efficient semantic communications. TokCom is motivated by the recent advent of powerful GFM/MLLMs based on transformer neural networks (NNs), where tokens are the basic processing units of text, images, audio/video signals that may represent words, image patches, temporal audio slices, or video subframes. GFMs/MLLMs-based token processing enables to encode the semantic content of multimodal signals at the transmitter and recover the corresponding semantics at the receiver leveraging the cross-modal context information, as depicted in Fig. 1. If a corrupted packet leads to an incomplete received message like “A beach with palm [MASK] and clear blue water”, where the missing word is replaced by a [MASK] token, TokCom can leverage a pre-trained LLM to predict the masked word based on the surrounding context, outputting “trees” to complete the sentence, thereby avoiding packet re-transmissions
@@
  • We introduce a novel TokCom framework along with its basic designs to leverage cross-modal context information in the semantic source compression, semantic channel coding, semantic multiple access, and semantic networking setups. TokCom integrates transformer-based next/masked token prediction via GFM/MLLMs.
  • We introduce and demonstrate a token-level loss/error mitigation scheme that leverages token likelihood estimation based on cross-modal context to predict and mitigate the tokens corrupted or lost in communication. In a typical image semantic communication setup, we demonstrate a significant improvement of the bandwidth efficiency, by TokCom leveraging the context information.
@@ II Tokenization, and Embedding of Various Data Modalities
To process large multimodal data, it is first segmented into various chunks, each consisting of several tokens. Tokens are then assigned unique IDs. Each token ID is mapped to a vector that represents the semantic meaning of the token in a dense, fixed-dimensional space, i.e., the embedding space. The embedding vectors are learned via pre-training to capture both syntactic and semantic relationships between tokens. These embeddings allow multimodal language models process inputs more effectively by capturing token semantics as well as the relationships between different tokens, i.e., context. A Tokenizer generates a large unified codebook of token embeddings that represent the whole corpus of multimodal data, e.g., colossal clean crawled corpus (C4) for text, LAION-5B for image-text pairs, and then the (pre-)trained MLLM essentially learns relations between tokens capturing the semantics and context. As an example, a vision language model, e.g., the contrastive language-image pretraining (CLIP) model [^1] https://github.com/openai/CLIP, learns through pre-training on large corpus of image-text pairs to relate the tokens of a “book” image with its corresponding text token. Similarly, the CLAP [^2] https://github.com/LAION-AI/CLAP and AudioCLIP [^3] https://github.com/AndreyGuzhov/AudioCLIP models learn the cross-modal relations via a shared embedding space for Text-Audio and Text-Image-Audio modalities.
@@
Text Tokenization/Embedding: In text, a token can represent a word, part of a word, or even a single character, depending on the tokenization model. Tokenization algorithms like WordPiece, byte-pair encoding (BPE), and unigram language modeling split words into subwords, with each subword mapped to a unique token in the vocabulary. Tokens are then projected into a high-dimensional feature space using word embeddings. Early models like Word2Vec and GloVe used fixed embeddings, while modern transformer models such as bidirectional encoder representations from transformers (BERT) and GPT learn contextual embeddings, meaning the representation of a word like “bank” varies depending on its context.
@@
Image/Video Tokenization/Embedding: Image tokens are created by dividing the image into fixed-size patches, which are then flattened and embedded. The challenge lies in balancing image fidelity and patch size; smaller patches can increase computational load, while larger ones may lose detail. For instance, vision transformer (ViT) uses $16\times 16$ pixel patches. After patchification and embedding, vector quatization is typically used to achieve a discrete latent space, where each image is represented by a sequence of token indices from a codebook.
@@
Audio Tokenization/Embedding: Audio signals are typically tokenized by transforming the waveform into spectrograms, typically using techniques such as log Mel filterbanks, which capture both the temporal and frequency-based features of the audio signal. The spectrograms are then divided into overlapping patches, and then flattened and projected into the embedding space to be processed by attention-based transformer models.
Finally, some commonly used tokenizers/models are BPE and WordPiece for text, VQ-VAE/GAN, TiTok[5] for image, VideoMAE [^4] https://github.com/MCG-NJU/VideoMAE and VidTok [^5] https://github.com/microsoft/VidTok for video, and HuBERT [^6] https://github.com/bshall/hubert for Audio signal modalities.
@@ III Tokens, Attention, and Transformers
Tokens are commonly used in various vision and natural language processing (NLP) models to represent data in a structured, numerical format that models can process. Transformer-based models, e.g., ViT, BERT, and GPTs, revolutionized token processing by introducing a fundamentally different approach to handling token sequences based on the attention mechanism. Transformers use self-attention mechanisms, where each token directly attends to all other tokens in the sequence, making it much easier for the model to capture long-range dependencies and contextual information. Compared to other NLP models, Transformers scale much more effectively [6], making it possible to train very deep models, such as the GPTs, LLaMA, or PaLM with billions of parameters. Transformer-based models may learn and predict contextual information either in a unidirectional or bidirectional fashion.
Unidirectional context learning/prediction refers to models that process input data in a single direction, e.g., left-to-right for text or row-by-row for images, where each token is predicted based only on the preceding tokens. Auto-regressive models like GPTs use this approach and are suited for tasks such as text completion or machine translation, where tokens are generated sequentially without access to future tokens, as depicted in Fig. 2(a).
Bidirectional context learning/prediction leverages the concept of masking, where random tokens in the input text or image are hidden, and the model learns to predict them by attending to both preceding and succeeding tokens. This bidirectional attention helps capture richer context. Models like BERT and MaskGIT [7] use masked modeling predict masked tokens by considering both nearby and distant tokens, as depicted in Fig. 2(b). Masked modeling is widely applied to generation tasks, such as image generation [7], where the model iteratively generates image tokens, revealing a portion of tokens at each step and progressively uncovering more with each iteration.
IV Token Communications: Opportunities and Challenges
TokCom provides the following main opportunities/advantages:
  • Tokens unify modalities and TokCom can leverages cross-modal relations to achieve ultra-low-bitrate SemCom. Furthermore, the in-context learning capabilities of MLLMs enables developing efficient TokCom frameworks that can adapt to various tasks, such as reconstruction, generation, and segmentation, among different modalities.
Table I provides a comparison between TokCom and the conventional SemCom schemes. Despite the above benefits, the large computational complexity of many existing GFMs/MLLMs is a challenge to be addressed in design of efficient TokCom systems. An effective solution to this would be collaborative cloud-edge-device TokCom based on task offloading, which we further investigate in the next sections.
@@ IV Token Communications: Opportunities and Challenges
Computational Complexity Typically uses smaller task-specific NN models. Typically based on computationally complex GFMs/MLLMs, GPTs, etc.
@@ V-A TokCom and Semantic Source Compression
TokCom systems can leverage GFMs/MLLMs-based context processing to achieve ultra-low-bitrate semantic source compression in future wireless networks. Language modeling is inherently a form of compression because it involves accurate prediction of the next/masked tokens in a sequence, thereby reducing uncertainty about the data. State-of-the-art GFM/MLLMs predict the conditional probabilities of tokens given context. The predictability-compression link is general across various data modalities, and by minimizing the log-likelihood loss during training, MLLMs implicitly optimize for compression. Thereby, compression is achieved when models effectively capture patterns and dependencies in the input multimodal data, making GFM/MLLMs inherently capable of functioning as efficient semantic compressors. As an example, the authors in [8] demonstrated that Chinchilla 70B, i.e., a language model primarily trained on text, can efficiently compress ImageNet patches to 43.4% of their original size and LibriSpeech audio samples to 16.4%, outperforming modality-specific compressors like PNG (58.5%) and FLAC (30.3%). Additionally, since tokenizers transform multimodal data into discrete tokens, tokenization inherently acts as a form of lossy compression. Its performance limits are dictated by the rate-distortion-perception theory [9], which defines the trade-off between rate, distortion, and the perceptual quality of the reconstructed signal. Moreover, MLLMs can selectively discard some of the tokens at the transmitter that are less relevant to downstream tasks, enabling token-based lossy compression. We demonstrate TokCom for semantic source compression in Fig. 3(a).
@@ V-B TokCom with Semantic Channel Coding
Another possibility is to leverage the semantic similarity of tokens to optimize a mapping from the token codebook to the channel symbols. For example, the two tokens “King” and “Queen” are semantically very similar. In fact, if the gender information is lost or mistaken due to channel errors, it can be simply inferred or corrected from numerous contextual clues at the receiver leveraging GFM/MLLMs. In other words, semantically similar tokens should be mapped onto channel symbols that are more probable to be confused during transmission. State-of-the-art multimodal embedding techniques enable the semantic similarity of tokens to be measured by their distance in the embedding space. We demonstrate this concept in Fig. 3(b).
Figure 3: Basic TokCom setups: (a) Semantic source compression: Left - lossless compression via arithmetic coding and MLLMs for probability prediction; right - discarding tokens less relevant to downstream tasks for token-based lossy compression. (b) Semantic channel coding: Map the embedding space to the modulation space; closer constellations for embedding vectors with high cosine similarity. (c) Semantic multiple access: Tokens from two users overlap in non-orthogonal multiple access, MLLMs separate sources using context and semantic orthogonality. (d) Semantic network protocols: Transmitter packetizes tokens and randomly drops for bitrate control; receiver predicts lost/errored packets using MLLMs.
@@ VI-A TokCom Setup
Each image is first tokenized into a sequence of $N=256$ discrete tokens. To facilitate efficient transmission, these tokens are grouped into packets, with each packet containing 4 tokens. To mitigate the impact of burst errors—where consecutive erroneous tokens can significantly degrade reconstruction quality due to the loss of contextual information—the token positions included within each packet are randomized. The token codebook size is $Q=1024$ and each token is represented by 10 bits, achieving the ultra-low-bitrate transmission of 0.039 bit per pixel (bpp).
For reliable transmission, each packet undergoes rate-$\frac{1}{2}$ convolutional encoding with cyclic redundancy check to enhance error detection and correction. The 16-QAM modulation is used and the bandwidth is 0.05 MHz. On the receiver side, demodulation and decoding are performed using a Soft Viterbi decoder, resulting in different packet error rate (PER) at different signal-to-noise ratio (SNR) values. Then, the average number of packet retransmissions required for successful data reception, denoted as $T$, is calculated as $\frac{1}{1-\text{PER}}$. Finally, we report the token-level performance metrics, token error rate (TER) and token communication efficiency defined as $\mathrm{TCE}=\frac{h\times w}{T\times N\times\log_{2}({Q})}$.